13) A 14-cm-diameter circular loop of wire is placed in a 1.48-T magnetic field.

A.When the plane of the loop is perpendicular to the field lines, what is the magnetic flux through the loop?

B.The plane of the loop is rotated until it makes a 38 ∘ angle with the field lines. What is the angle in the equation ΦB = BAcosθ for this situation?

C.What is the magnetic flux through the loop at this angle?
Answer #1

ϕ=

π  ⋅ (0.07m)^2×1•48T cos0°=7.25×10^-3π=22.78wb

Now for the second case

θ=90o−38°=52°

ϕ=π  ⋅ (0.07m)2 1.48T  cos62°=10.69wb